OpenStudy (anonymous):

Solve. sqrt6^(8x)=216^(x-3)

OpenStudy (lgbasallote):

(6^1/2)^(8x) = (6^3)^(x-3) 6^(4x) = 6^(3x - 9) 4x = 3x - 9 x = -9